# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v2

@@ 3.5.3 PIM/NDP
Computing Enhancement. (1) PIM-NPU. NeuPIMs [197] and IANUS [198] are heterogeneous PIM acceleration systems that jointly exploits a conventional GEMM-focused NPU and GEMV-optimized PIM devices. NeuPIMs [197] first proposes dual row buffers in each bank, facilitating the simultaneous management of memory read/write operations and PIM commands, to enable concurrent operations on both NPU and PIM platforms. Further, it employs a runtime sub-batch interleaving technique to maximize concurrent execution for the inherent dependencies between GEMM and GEMV in LLMs. Our evaluation demonstrates that NeuPIMs with $\sim$76W (0.6348W+75W) power consumption (32GB HBM) achieves $\sim$3k tokens/s with batch size 8 for GPT3-7B. IANUS [198] proposes novel PIM access scheduling that manages not only the scheduling of normal memory accesses and PIM computations but also workload mapping across the PIM and the NPU. The evaluations show that for GPT-6.7B model, IANUS achieves 127.1 tokens/s with about 240W power consumption. Cambricon-LLM [209] proposes a chiplet-based hybrid architecture with NPU and a dedicated NAND flash chip to enable efficient LLM inference. It utilizes both the high computing capability of NPU and the data capacity of the NAND flash chip. The NAND flash chip is enhanced by in-flash computing to perform precise lightweight on-die processing, and the NPU performs matrix operations and handles special function computations. Experimental results show that Cambricon-LLM with $\sim$37W power consumption achieves 3.44 tokens/s ($\sim$) and 36.34 tokens/s for 70B and 7B LLMs, which is over 22$\times$ to 45$\times$ faster than existing flash-offloading technologies, respectively. (2) PIM-GPU. MoNDE [199] is a near-data computing solution that efficiently enables Mixture-of-Experts (MoE) LLM inference on heterogeneous hardwares. MoNDE reduces the volume of MoE parameter movement by transferring only the hot experts to the GPU, while computing the remaining cold experts inside the host memory device. MoNDE can achieve inference latency comparable to an ideal GPU system with infinite memory. AttAcc [200, 201] is also an heterogeneous system equipped with GPUs to accelerate the attention layers. Compared to the monolithic state-of-the-art GPU system, AttAcc achieves significantly higher throughput (up to 2.81$\times$) and energy efficiency (up to 2.67$\times$) for GPT3-175B model. In 2024, SK Hynix proposes LPDDR6-based heterogeneous LLM system called AiMX-xPU[208], which consists 1 NVIDIA H100 GPU and 3 AiMX, achieving 167 tokens/s, 220 tokens/s, and 900 tokens/s with batch size 1, 8, and 32 for OPT-30B. (3) PIM-FPGA. Kang et al. [202] put the memory-bound GEMV calculations including projections, feed-forward-network layer, and the last fully-connected layer with pre-trained weight on HBM2-PIM and evaluate system with PIM-powered Xilinx Alveo U280 board. For GPT-1.3B model, after scaling the bandwidth by 8$\times$ for PIM technology, it can achieve about 347.83 tokens/s, which is 1.6$\times$ faster than the NVIDIA A100 GPU. From 2022 to 2023, SK Hynix proposes GDDR6-based accelerator-in-memory named AiM [228, 229] and integrates it with 2 XLINX XCVU9P FPGA chips for control and communication as a prototype called AiMX [207] to achieve 330 tokens/s for OPT-6.7B with batch size 1. (4) PIM-TPU. H3D-Transformer [204] proposes a heterogeneous 3D-based accelerator design for transformer models, which adopts an interposer substrate with multiple 3D memory/logic hybrid cubes optimized for accelerating different MatMul workloads. An approximate computing scheme is proposed to take advantage of heterogeneous computing paradigms of mixed-signal compute-in-memory (CIM) and digital tensor processing units (TPU). From the system-level evaluation results, 10 TOPS/W energy efficiency is achieved for the GPT2 model, which is about 2.6$\times$-3.1$\times$ higher than the baseline with 7nm TPU and stacked FeFET memory. 3D-HI [206] leverage chiplet-based heterogeneous integration to design a Network-on-Interposer (NoI) architecture to accelerate LLM inference. 3D-HI uses the streaming multiprocessors along with the associated memory controllers (SM-MCs) for multi-head attention and ReRAM cores for the feed-forward network, which optimize both achievable energy efficiency and throughput. Further, vertical integration on top of a 2.5D interposer helps to enhance overall system performance and alleviates the issue of memory bottlenecks. Experimental results demonstrate that 3D-HI lowers the latency and energy consumption by up to 22.8$\times$ and 5.36$\times$ with respect to an equivalent state-of-the-art chiplet-based platform.
@@
Figure 12: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with heterogeneous cooperation.
@@
Memory Enhancement. (1) PIM-CXL. As the frequent transfers of these model parameters and intermediate values are performed over relatively slow device-to-device interconnects such as PCIe or NVLink, they become the key bottleneck for efficient acceleration of LLMs. Kim et al. [203] exploit PIM, which uses bank-level parallelization to provide higher internal memory bandwidth compared to traditional DRAM, resulting in a significant increase in on-DRAM compute bandwidth. In addition to achieving high capacity through Compute eXpress Link (CXL) memory expansion, CXL-PNM demonstrates performance improvements by integrating computational logic into memory products, consequently increasing internal memory bandwidth. Evaluation results show that the performance of memory-bounded LLMs was significantly improved with PIM and PNM by up to 4.5$\times$ and 4.4$\times$, respectively. CXL-PNM [205] first devises an LPDDR5X-based CXL memory architecture with 512GB of capacity and 1.1TB/s of bandwidth, which boasts 16$\times$ larger capacity and 10$\times$ higher bandwidth than GDDR6 and DDR5-based CXL memory architectures, respectively, under a module form-factor constraint. Second, it designs a CXL-PNM controller architecture integrated with an LLM inference accelerator, exploiting the unique capabilities of such CXL memory to overcome the disadvantages of competing technologies such as HBM-PIM and AxDIMM. The evaluation shows that for OPT-13B model, CXL-PNM achieves 42.68 tokens/s with 77.6W power consumption. (2) PIM-3D Stack. Sharda et al. [129] propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption.
3.5.4 Comparison
@@
In Figure 12, for heterogeneous cooperation, power consumption ranges from 6.3W to 575W, with inference speeds between 4.9 tokens/s and 330 tokens/s. The energy efficiency ranges from 0.0145 tokens/J to 4.342 tokens/J. PowerInfer-2 [196] with Snapdragon 8 Gen 3 (CPU) achieves the lowest power consumption with 6.3W. SK Hynix AiMX [207] (PIM/NDP) achieves the highest throughput with batch size 1 and the highest energy efficiency.
@@
3.6 Homogeneous Cooperation
3.6.1 Overview
@@
3.6.2 CPU
Distributed computing emerges as a prevalent strategy to mitigate single-node memory constraints and expedite LLM inference performance. He et al. [210] propose an efficient distributed inference optimization solution for LLMs on CPUs. On four 5th Gen Intel Xeon Scalable Processors the result shows the generation speed on Qwen-72B is 7.14 tokens/s. He et al. [211] also propose new attention flow, SlimAttention, to reduce the KV cache size and ensure precision for efficient LLM inference on CPUs. The experimental results on Llama2-70b shows the latency of token generation is 4 tokens/s with 2 sockets and 11.4 tokens/s with 8 sockets on Intel Xeon 8563C CPUs.
3.6.3 FPGA
DFX [212] is a multi-FPGA acceleration which uses model parallelism and optimized dataflow to improve LLM inference speed in both prefilling and decoding phases. With the implementation on four Xilinx Alveo U280 FPGAs, DFX achieves about 120 tokens/s for GPT2-1.5B model.
4 Discussion
4.1 Setup
@@
  • For most ASIC platforms, we first define a post-silicon work [126] with an absolute inference speed as the baseline and scale the results by the peak computational performance (TOPS or GOPS).
@@
4.2 Hardware Comparison
4.2.1 Small batch size (bs=1)
@@
Figure 13: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with different optimization methods.
Figure 14: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 8) vs power on different platforms with different optimization methods.
4.2.2 Large batch size (bs=8)
@@
4.3 Optimization Method Comparison
4.3.1 Small batch size (bs=1)
@@
4.3.2 Large batch size (bs=8)
@@
In Figure 14, on GPU platforms, the methods used include quantization, sparsity, and operator optimization. Quantization and operator optimization achieve higher absolute inference speeds, reaching approximately 1000 tokens/s and 1033 tokens/s, respectively, while sparsity achieves about 290 tokens/s. Regarding efficiency, quantization, and operator optimization also achieve higher energy efficiency with 3.155 tokens/J and 2.3275 tokens/s compared to sparsity (0.725 tokens/J). On FPGA platforms, the methods employed include quantization and sparsity, which are often used together. The highest inference speed reaches 370 tokens/s, with an efficiency of up to 2.056 tokens/J. On ASIC platforms, the methods employed include operator optimization. It can achieve higher absolute inference speeds, reaching approximately 1176.47 tokens/s with 7.843 tokens/J energy efficiency. On PIM/NDP platforms, the main methods used are heterogeneous cooperation. Heterogeneous cooperation can achieve an inference speed of about 3000 tokens/s and 40 tokens/J.
Figure 15: LLM ($\sim$ 7 billion parameters) inference for different edge-side scenarios.
4.4 Discussion on Edge-side Scenarios
In Figure 15, we also analyze two typical edge-side LLM inference application scenarios: AI smartphones and robots. For the AI smartphones, edge hardware typically requires power consumption to be limited to 10W or less, and with human reading speeds around 10 tokens/s [230], the inference speed requirement is relatively low. The lower left corner of the plot corresponds to this scenario. A notable example is the Qualcomm Snapdragon 8 Gen 3 chip [65], which utilizes quantization techniques to optimize the model, allowing it to achieve the necessary inference speed while staying under the 10W power limit. And MARCA [183] and Tender [126] can also satisfy the inference demand. Such optimization strategies have enabled some edge CPUs and ASICs to meet the inference needs of smartphone applications. For the robots, the requirements for inference speed are significantly higher. To support real-time systems with an operational frequency of 100-1000Hz [231], the inference speed must reach 100-1000 tokens/s, while the hardware power consumption typically needs to reach around 20W. According to the data in the plot, although some pre-silicon designs meet this requirement, they are currently limited to the pre-simulation stage and have not yet undergone post-silicon validation. For instance, Guo et al. [127] and SpecPIM [168] show promising results in simulation, indicating that they could meet the power and speed demands, but whether they can actually fulfill the requirements for robot applications still requires further post-silicon validation and tape-out verification.
5 Conclusion

### v3

@@ 3.5.3. PIM/NDP
Computing Enhancement. (1) PIM-NPU. NeuPIMs (heo2024neupims, ) and IANUS (seo2024ianus, ) are heterogeneous PIM acceleration systems that jointly exploits a conventional GEMM-focused NPU and GEMV-optimized PIM devices. NeuPIMs (heo2024neupims, ) first proposes dual row buffers in each bank, facilitating the simultaneous management of memory read/write operations and PIM commands, to enable concurrent operations on both NPU and PIM platforms. Further, it employs a runtime sub-batch interleaving technique to maximize concurrent execution for the inherent dependencies between GEMM and GEMV in LLMs. Our evaluation demonstrates that NeuPIMs with $\sim$76W (0.6348W+75W) power consumption (32GB HBM) achieves $\sim$3k tokens/s with batch size 8 for GPT3-7B. IANUS (seo2024ianus, ) proposes novel PIM access scheduling that manages not only the scheduling of normal memory accesses and PIM computations but also workload mapping across the PIM and the NPU. The evaluations show that for GPT-6.7B model, IANUS achieves 127.1 tokens/s with about 240W power consumption. Cambricon-LLM (yu2024cambricon, ) proposes a chiplet-based hybrid architecture with NPU and a dedicated NAND flash chip to enable efficient LLM inference. It utilizes both the high computing capability of NPU and the data capacity of the NAND flash chip. The NAND flash chip is enhanced by in-flash computing to perform precise lightweight on-die processing, and the NPU performs matrix operations and handles special function computations. Experimental results show that Cambricon-LLM with $\sim$37W power consumption achieves 3.44 tokens/s ($\sim$) and 36.34 tokens/s for 70B and 7B LLMs, which is over 22$\times$ to 45$\times$ faster than existing flash-offloading technologies, respectively. (2) PIM-GPU. MoNDE (kim2024monde, ) is a near-data computing solution that efficiently enables Mixture-of-Experts (MoE) LLM inference on heterogeneous hardwares. MoNDE reduces the volume of MoE parameter movement by transferring only the hot experts to the GPU, while computing the remaining cold experts inside the host memory device. MoNDE can achieve inference latency comparable to an ideal GPU system with infinite memory. AttAcc (choi2023unleashing, ; park2024attacc, ) is also an heterogeneous system equipped with GPUs to accelerate the attention layers. Compared to the monolithic state-of-the-art GPU system, AttAcc achieves significantly higher throughput (up to 2.81$\times$) and energy efficiency (up to 2.67$\times$) for GPT3-175B model. In 2024, SK Hynix proposes LPDDR6-based heterogeneous LLM system called AiMX-xPU(kim2024sk, ), which consists 1 NVIDIA H100 GPU and 3 AiMX, achieving 167 tokens/s, 220 tokens/s, and 900 tokens/s with batch size 1, 8, and 32 for OPT-30B. (3) PIM-FPGA. Kang et al. (kang2023era, ) put the memory-bound GEMV calculations including projections, feed-forward-network layer, and the last fully-connected layer with pre-trained weight on HBM2-PIM and evaluate system with PIM-powered Xilinx Alveo U280 board. For GPT-1.3B model, after scaling the bandwidth by 8$\times$ for PIM technology, it can achieve about 347.83 tokens/s, which is 1.6$\times$ faster than the NVIDIA A100 GPU. From 2022 to 2023, SK Hynix proposes GDDR6-based accelerator-in-memory named AiM (lee20221ynm, ; kwon2022system, ) and integrates it with 2 XLINX XCVU9P FPGA chips for control and communication as a prototype called AiMX (AiMX, ) to achieve 330 tokens/s for OPT-6.7B with batch size 1. (4) PIM-TPU. H3D-Transformer (luo2024h3d, ) proposes a heterogeneous 3D-based accelerator design for transformer models, which adopts an interposer substrate with multiple 3D memory/logic hybrid cubes optimized for accelerating different MatMul workloads. An approximate computing scheme is proposed to take advantage of heterogeneous computing paradigms of mixed-signal compute-in-memory (CIM) and digital tensor processing units (TPU). From the system-level evaluation results, 10 TOPS/W energy efficiency is achieved for the GPT2 model, which is about 2.6$\times$-3.1$\times$ higher than the baseline with 7nm TPU and stacked FeFET memory. 3D-HI (sharma2023heterogeneous, ) leverage chiplet-based heterogeneous integration to design a Network-on-Interposer (NoI) architecture to accelerate LLM inference. 3D-HI uses the streaming multiprocessors along with the associated memory controllers (SM-MCs) for multi-head attention and ReRAM cores for the feed-forward network, which optimize both achievable energy efficiency and throughput. Further, vertical integration on top of a 2.5D interposer helps to enhance overall system performance and alleviates the issue of memory bottlenecks. Experimental results demonstrate that 3D-HI lowers the latency and energy consumption by up to 22.8$\times$ and 5.36$\times$ with respect to an equivalent state-of-the-art chiplet-based platform.
@@
Figure 12. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with heterogeneous cooperation.
@@
Memory Enhancement. (1) PIM-CXL. As the frequent transfers of these model parameters and intermediate values are performed over relatively slow device-to-device interconnects such as PCIe or NVLink, they become the key bottleneck for efficient acceleration of LLMs. Kim et al. (kim2024breakthrough, ) exploit PIM, which uses bank-level parallelization to provide higher internal memory bandwidth compared to traditional DRAM, resulting in a significant increase in on-DRAM compute bandwidth. In addition to achieving high capacity through Compute eXpress Link (CXL) memory expansion, CXL-PNM demonstrates performance improvements by integrating computational logic into memory products, consequently increasing internal memory bandwidth. Evaluation results show that the performance of memory-bounded LLMs was significantly improved with PIM and PNM by up to 4.5$\times$ and 4.4$\times$, respectively. CXL-PNM (park2024lpddr, ) first devises an LPDDR5X-based CXL memory architecture with 512GB of capacity and 1.1TB/s of bandwidth, which boasts 16$\times$ larger capacity and 10$\times$ higher bandwidth than GDDR6 and DDR5-based CXL memory architectures, respectively, under a module form-factor constraint. Second, it designs a CXL-PNM controller architecture integrated with an LLM inference accelerator, exploiting the unique capabilities of such CXL memory to overcome the disadvantages of competing technologies such as HBM-PIM and AxDIMM. The evaluation shows that for OPT-13B model, CXL-PNM achieves 42.68 tokens/s with 77.6W power consumption. (2) PIM-3D Stack. Sharda et al. (sharda2024accelerator, ) propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption.
3.5.4. Comparison
@@
In Figure 12, for heterogeneous cooperation, power consumption ranges from 6.3W to 575W, with inference speeds between 4.9 tokens/s and 330 tokens/s. The energy efficiency ranges from 0.0145 tokens/J to 4.342 tokens/J. PowerInfer-2 (xue2024powerinfer, ) with Snapdragon 8 Gen 3 (CPU) achieves the lowest power consumption with 6.3W. SK Hynix AiMX (AiMX, ) (PIM/NDP) achieves the highest throughput with batch size 1 and the highest energy efficiency.
@@
3.6. Homogeneous Cooperation
3.6.1. Overview
@@
3.6.2. CPU
Distributed computing emerges as a prevalent strategy to mitigate single-node memory constraints and expedite LLM inference performance. He et al. (he2024distributed, ) propose an efficient distributed inference optimization solution for LLMs on CPUs. On four 5th Gen Intel Xeon Scalable Processors the result shows the generation speed on Qwen-72B is 7.14 tokens/s. He et al. (he2024inference, ) also propose new attention flow, SlimAttention, to reduce the KV cache size and ensure precision for efficient LLM inference on CPUs. The experimental results on Llama2-70b shows the latency of token generation is 4 tokens/s with 2 sockets and 11.4 tokens/s with 8 sockets on Intel Xeon 8563C CPUs.
3.6.3. FPGA
DFX (hong2022dfx, ) is a multi-FPGA acceleration which uses model parallelism and optimized dataflow to improve LLM inference speed in both prefilling and decoding phases. With the implementation on four Xilinx Alveo U280 FPGAs, DFX achieves about 120 tokens/s for GPT2-1.5B model.
4. Further Comparison
4.1. Setup
@@
  • For most ASIC platforms, we first define a post-silicon work (10609625, ) with an absolute inference speed as the baseline and scale the results by the peak computational performance (TOPS or GOPS).
@@
4.2. Hardware Comparison
4.2.1. Small batch size (bs=1)
@@
Figure 13. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with different optimization methods.
Figure 14. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 8) vs power on different platforms with different optimization methods.
4.2.2. Large batch size (bs=8)
@@
4.3. Optimization Method Comparison
4.3.1. Small batch size (bs=1)
@@
4.3.2. Large batch size (bs=8)
@@
In Figure 14, on GPU platforms, the methods used include quantization, sparsity, and operator optimization. Quantization and operator optimization achieve higher absolute inference speeds, reaching approximately 1000 tokens/s and 1033 tokens/s, respectively, while sparsity achieves about 290 tokens/s. Regarding efficiency, quantization, and operator optimization also achieve higher energy efficiency with 3.155 tokens/J and 2.328 tokens/s compared to sparsity (0.725 tokens/J). On FPGA platforms, the methods employed include quantization and sparsity, which are often used together. The highest inference speed reaches 370 tokens/s, with an efficiency of up to 2.056 tokens/J. On ASIC platforms, the methods employed include operator optimization. It can achieve higher absolute inference speeds, reaching approximately 1176.47 tokens/s with 7.843 tokens/J energy efficiency. On PIM/NDP platforms, the main methods used are heterogeneous cooperation. Heterogeneous cooperation can achieve an inference speed of about 3000 tokens/s and 40 tokens/J.
5. Further Discussion
The development of edge intelligence has gained significant momentum, driven by the increasing capability of LLMs and the increasing demands of edge applications. As we look toward the future, three trends emerge: multimodality, inference-time compute, and higher inference energy efficiency, which is promising to redefine the capabilities of edge AI systems, as shown in Fig. 15. Multimodality focuses on integrating diverse data types, such as visual, auditory, textual, and sensory inputs, to enhance the system’s ability equipped with LLMs to understand and adapt to complex scenarios, supporting a wide range of edge devices like PCs, smartphones, AR glasses, and smartwatches. Inference-time compute emphasizes the deep thinking during LLM inference leveraging iterative processes such as Chain-of-Thought (CoT) reasoning to continuously improve model capabilities. It is achieved by performing a post-training process. Higher energy efficiency highlights the demand for increasing energy efficiency from 1 tokens/J to >10 tokens/J for high real-time scenarios, such as embodied intelligence represented by robotics control and autonomous driving.
Figure 15. The trends of edge AI system: multimodality, inference-time computation, and higher inference energy efficiency.
5.1. Multimodality
As shown in Fig. 16, we summarize the release dates and the number of training text tokens for typical LLMs over the past five years. The number of training tokens has been growing at a rate of 2.5$\times$ per year. GPT-3 is trained using only 0.4 trillion text tokens, while Qwen2.5 and Llama3.1 use 18 trillion and 15 trillion tokens, respectively. According to the result of Epoch AI, the total number of human-generated public text tokens is $\sim$300 trillion (villalobos2024run, ). At this growth rate, it is expected that by 2027, we will face a shortage of available public text tokens for training LLMs. In addition to text tokens, there is a vast amount of other modalities of tokens on the Internet, such as image and video tokens (total $\sim$1650 trillion (villalobos2024run, )). Therefore, from the perspective of token availability, by leveraging multiple modalities of tokens for training LLMs, we can continue the LLMs’ scaling law (kaplan2020scaling, ). From the perspective of human interaction, humans rely on various senses as gateways to interact with the physical world, which is inherently multimodal. Recent research has also proven that multimodal learning provably performs better than single modal due to the more accurate estimate of the latent space representation (huang2021makes, ). These factors indicate that future LLMs should have the capability of multimodal interactions to perceive, process, and generate information across multiple modalities.
Figure 16. The number of training text tokens for typical single modal LLMs in the past 5 years. We will lack the public text tokens after 2027. By using multimodal tokens for training, we can continue the LLMs’ scaling law.
Currently, more than 80% multimodal LLMs focus on both vision and language tasks (zhang2024mm, ). As shown in Fig. 17, their development can be roughly divided into three stages. First, from May to September 2023, the model size is small and the accuracy ranges from 20% to 40%. Then, from December 2023 to April 2024, with the release of GPT4v, the model accuracy significantly improves, reaching nearly 60%. After May 2024 with the release of GPT4o, more multimodal models are released and the accuracy reaches almost 80%. It should be noted that the development of multimodal LLMs (MLLMs) heavily relies on the construction of multimodal datasets. And another significant trend in MLLMs is the concept of native multimodality. By aligning diverse modalities during the training stage, native MLLMs enable end-to-end inputs and outputs.
Figure 17. The timeline of multimodal (Vision-Language) LLMs: average accuracy, publication date, and model size. The results are collected from OpenCompass (OpenCompass, ).
5.2. Inference-time Compute
In the LLM era, data is the fossil fuel, while the age of pre-training will end because the data is not growing (llyanips, ). Inference-time compute, as represented by OpenAI’s o1 paradigm (openai_o1, ), is now seen as the critical pathway to enhance model capabilities in the next LLM era. The principles of o1 can be illustrated through its distinction from GPT4, as shown in Fig. 18(a). During the training stage, GPT4 follows the conventional workflow: pre-training a base model followed by fine-tuning through reinforcement learning from human feedback (RLHF). In contrast, o1 introduces several innovations in the post-training stage. Beyond incorporating reinforcement learning (RL), o1 places particular emphasis on safety optimization to ensure that o1 is equipped for more nuanced tasks and safe decision-making during deployment. The difference between GPT4 and o1 becomes more pronounced in the inference stage. GPT4 generates answers directly without explicitly involving iterative reasoning. By comparison, o1 introduces a “Think&Summary” phase before producing answers. This step leverages advanced techniques such as Chain-of-Thought (CoT) reasoning and tree search to perform more intricate logical reasoning and step-by-step problem solving. Compared to GPT4o, the o1 paradigm achieves a 50% improvement in accuracy on the complex logical reasoning tasks like AI Math Evaluation (AIME) test dataset. However, the inference cost is increased by 10-100$\times$ (o1inferencecost, ).
Inference-time compute can enhance the model capabilities but also substantially increase inference time, particularly for complex reasoning tasks. As shown in Fig. 18(b), in the GSM8K benchmark, we employ Llama2-7B as the backbone LLM, combined with the open-source LLM-Reasoners (hao2024llm, ). The inference time is increased by 678$\times$ compared to using Llama2-7B directly. This substantial time overhead primarily stems from the iterative multi-step process inherent in inference-time compute, which is necessary to satisfy the requirements of more complex reasoning task. Therefore, choosing the appropriate scenarios and conditions for employing inference-time compute remains a topic worthy of further research. Additionally, inference-time compute also changes the runtime breakdown of the LLM inference. As shown in Fig. 18(c), the proportion of the prefill stage significantly increases from 1.5% to 23.5%, while the decode stage decreases from 98.5% to 54.8%. The increase in the prefill stage is partly due to the need to merge the human input tokens with the template input tokens as the whole LLM input, which effectively extends the input sequence length. Moreover, in the iterative process, intermediate output tokens are cumulatively fed back as inputs, leading to a gradual increase in the sequence length for prefill. Meanwhile, the introduction of the Process Reward Model (PRM) accounts for 21.7% of the time. This shift of runtime breakdown highlights that future optimization methods for inference-time compute should consider all prefill stage, decode stage, and PRM simultaneously.
Inference-time compute is poised to become the foundational LLM inference paradigm of next-generation AI systems. For edge applications like autonomous driving and robotics, the advanced model capabilities are essential while the exponential growth of inference cost presents a significant challenge for both hardware and software. Real-time responsiveness, higher throughput, and efficient resource utilization are becoming essential attributes for supporting these edge applications.
Figure 18. Inference-time Compute: (a) the difference of training and inference stages, (b) the increment of inference time, and (c) runtime breakdown of inference time.
5.3. Higher Inference Energy Efficiency
In future edge applications, particularly in high real-time scenarios such as embodied intelligence represented by robotics control and autonomous driving, the system’s control frequency needs to reach 100-1000 Hz (stewart1990implementing, ; bena2023hybrid, ; mafi2009parallel, ; atkins2023mintnet, ). This implies that systems using LLMs as the decision-making center for actions or instructions should achieve an inference speed of at least 100-1000 tokens/s. Since edge devices need to consider energy supply, the power of edge chips should be <20W (the power consumption of human brain (markram2012human, ; magistretti2015cellular, )) and then the inference efficiency should exceed 10 tokens/J. However, as shown in the Fig. 19, current commercial edge chips such as Tesla RPU (teslarpu, ) and NVIDIA Jetson Orin (nvidiaorin, ), even with software optimization methods, only achieve an actual inference speed of 10-30 tokens/s and an efficiency of <1 tokens/J. This presents a gap of 1-2 orders of magnitude in both absolute inference speed and efficiency compared to future edge application demands.
To achieve this goal, we should continue the algorithm-hardware co-design approach. On the algorithm side, more aggressive model compression methods can be employed with high-quality datasets for retraining to minimize accuracy loss. On the hardware side, in addition to designing specific hardware units to support algorithm optimizations, it is necessary to design more innovative chip architectures, such as 3D DRAM-stacked integrated architectures that significantly reduce data transfer latency and energy consumption, dataflow architectures that optimize data transfer paths to improve computational efficiency, and wafer-scale architectures that provide higher computational and storage capabilities through ultra-large-scale integration. These co-design approaches can effectively bridge the current gaps in inference speed and energy efficiency of edge-side chips, meeting the demands of future applications.
Figure 19. The performance gap between the future and the current edge AI systems.
6. Conclusion